help me solve this problem please,
Suppose you work in a lab, and you need a 33% acid solution. You have a 5% solution and a 40% solution in stock. You need 10 liters of the 33% acid solution. how many liters of 5% solution and 40% solution should you use?

Answer by ankor@dixie-net.com(22740)

Suppose you work in a lab, and you need a 33% acid solution.
You have a 5% solution and a 40% solution in stock.
You need 10 liters of the 33% acid solution.
how many liters of 5% solution and 40% solution should you use?
:
let a = amt of 40% solution
the resulting amt is to be 10 liters, therefore
(10-a) = amt of 5% solution
:
A typical mixture equation
40a + 5(10-a) = 33(10)
40a + 50 - 5a = 330
40a - 5a = 330 - 50
35a = 280
a = 280/35
a = 8 liters of the 40% solution
then
10 - 8 = 2 liters of the 5% solution